A square shaped garden has a diagonal that measures 15 feet what are the dimensions of the garden

Answer by checkley77(12569)

x^2+x^2=15^2
2x^2=225
x^2=225/2
x^2=112.5
x=sqrt/112.5
x=10.6 ans.